If Xsquare - 5x+1= 0 then find the the value of x power5 + 1/ x power5

If 
    Xsquare - 5x+1= 0 then find the the value of x power5 + 1/ x power5

Grade:10

2 Answers

Anand Chauhan
5 years ago
x2-5x+1=0
x2+1=5x
now on dividing both side by x we get
x+ 1/x = 5
now squaring both sides
x2+ 1/x2+2=25
x2+ 1/x2=23  let this be equation number 1
Now multiply it by x+ 1/x
(x2+1/x2)(x+ 1/x)= (23)(5)           ---------(since x+ 1/x = 5  (see third step ))
x3 + 1/x + (x+ 1/x) = 115
x3 + 1/x = 110 ---------(since x+ 1/x = 5  (see third step ))     (Now let this be equation number 2)
now multiply equation number 1 and 2
(x2+ 1/x2)(x3 + 1/x3) = (23)(110)
x5+1/x5+ (x+ 1/x) = 2530
x5+1/x5=2530-5
x5+1/x5=2525
Mona
3 years ago
Find the value of x when X2-3x+1=0 then X10+1÷x10 is find find find find find find find find find find find find